« Project Status: 2 January 2008 2030UTC | Main | Project Status: 4 January 2008 2030UTC »

Project Status: 3 January 2008 2030UTC

IOP #4 Underway Today

Weather Status: Rainbands at 2030UTC (1230L) were in the Central Valley west of the ARB. Extrapolation puts the leading edge of these showers in the ARB 22-23UTC (14-15L). Most intense rain in the first surge are forecast to reach the ARB by tonight and begin the accumulation of significant precipitation. Precipitation should continue through Friday as rain below BLU and snow above. A cold frontal passage is predicted by 5/00UTC (1600L) that should turn precipitation to snow above 3000ft. By Saturday morning we should see 3 to 6 inches of precipitation (coarser grid models on the low end; higher res models including the HMT ensemble favoring the higher amounts). Feet of snow possible at the elevation of BLU. Prior to the front, high winds are expected with gusts 50-70kts. After the front the winds subside but the heavy precipitation continues turning to snow at BLU by afternoon. There is some uncertainty about this changeover and it will be monitored closely. Snow will continue through Sunday morning with a break likely. Another upper level trough will approach the ARB Sunday night and provide another round of precipitation through early Tuesday. A warmer event is likely in the late Tues to Wed (8-9 Jan) time frame.

Summary: IOP #4 3 Jan/21UTC through 6 Jan/ 12UTC; 2nd event Sunday night to Tuesday; 3rd event Tuesday night to Wednesday.

Observational Status: Radar operations will begin a bit earlier with a start time of 3/22UTC or as soon as the crews can deploy to the site. RAOBs at SHS, RNO, and OAK will begin as planned at 3/21UTC. With warm advection generally ending by early Saturday RAOBS will end Saturday morning with the last release at 5/12UTC at all sites. With heavy snows likely at BLU, radar personnel may be on site for an extended period. If there is no clear Sunday break or likely closed highways near BLU (implying a 5-6 day stay) a decision to abandon the site may be necessary tomorrow. The Chico profiler site is again operative.

Decision: RAOBs 3 Jan/ 21 UTC to 5 Jan 12 UTC - 4hr intervals; Radar 3 Jan/ 21-22 UTC to 6 Jan/ 12 UTC (unless terminated early).

Product Status: HMT high resolution model runs were reestablished with output now reaching the HMT 2008 web site. Forecasts are generally supportive of the higher expected precipitation amounts. Cross-sections show rapidly falling freezing levels during the day on Friday, possibly resulting in earlier snow at BLU.

McGinley, GSD

TrackBack

TrackBack URL for this entry:
http://www.etl.noaa.gov/cgi-bin/mt/mt-tb.cgi/540